OpenStudy (anonymous):

Factor the following polynomial: x2 + 16x + 48 (x + 6)(x + 8) (x + 16)(x + 1) (x + 4)(x + 12) (x + 8)(x + 2)

OpenStudy (anonymous):

first one :)

OpenStudy (anonymous):

x^2 + 16x + 48 = x^2 + 4x +12x + 48 = x(x+4) + 12(x+4) = (x+12)(x+4)

OpenStudy (anonymous):

The first one is NOT correct. Try to expand it. (x + 6)(x + 8) = x(x + 8) + 6(x + 8) = x^2 + 8x + 6x + 48 = x^2 + 14x + 48